A train fare for a school child is half the fare of a teacher. The total fare for 120 children and 15teachers for an excursion is ghc180. C(1). Find the fare for a child. (2). How many children will go on the excursion with 20teachers for a total fare of ghc240.
Answer by Seutip(231)

Let
x - be the fare of the teacher
x/2 - be the fare of the child
120(x/2)+15(x)=180
60x+15x=180
75x=180
x= ghc 2.4; be the fare of the teacher
x/2= ghc 1.2; be the fare of the child

Let
y -be the number of child to go to the excursion with 20 teachers and a total fare of ghc 240
1.2(y)+2.4(20) = 240
1.2y+48 = 240
1.2y = 192
y = 160 students; will to go to the excursion with 20 teachers and a total fare of ghc 240
Thank you and stay safe!